OpenStudy (anonymous):

Separate 16 into two parts, one which of which shall be 3 times the other!

OpenStudy (anonymous):

Is this right? x = smaller number 3x = larger number x + 3x = 116 4x = 116 4x/4 = 116/4 x = 29

OpenStudy (anonymous):

cant be because the two numbers should add back up to 29

OpenStudy (anonymous):

NO they should add up to 116

OpenStudy (anonymous):

the number is 116 not 16 sorry lol

OpenStudy (anonymous):

oh okay that makes sense if the number is 116

OpenStudy (anonymous):

Can you help with my other ones please :D

OpenStudy (anonymous):

okay, ill try